Give exact and approximate solutions to three decimal places.
(x-9)^2 = 45

1) What are the exact solutions???
2) What are the approximate solutions???

Answer by stanbon(75887)

Give exact and approximate solutions to three decimal places.
(x-9)^2 = 45
1) What are the exact solutions???
(x-9)^2-45 = 0
[x-9-3sqrt(5))(x-9+3sqrt(5) = 0
----
x = 9+3sqrt(5) or x = 9-3sqrt(5)
--------------------------------------

2) What are the approximate solutions???
????
Cheers,
stan H.